Susan left her home, on foot, at 9:00 am at 1.5 mph. At 10:00 am her friend,
following the same route, ran at 5.25 mph to meet her. How many minutes, after
10:00 am, did the two friends meet?

Answer by palanisamy(496)

Let the two friends meet t minutes after 10.00 AM
In t minutes both friends have travelled equal distances
[(60+t)x1.5]/60 = [tx5.25]/60
90+1.5t = 5.25 t
90 = 3.75t
t = 90 / 3.75
t = 24 minutes